HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

C_{3}: RO-VIBRATIONAL ENERGY LEVELS AND LINE-STRENGTHS FOR THE ELECTRONIC GROUND-STATE

\underline{J. TENNYSON} , A.E. LYNAS-GRAY, S.N. YURCHENKO and O.L. POLYANSKY, Department of Physics and Astronomy, University College London,Gower Street, London WC1E 6BT; N.F. ZOBOV, Institute of Applied

Physics, Russian Academy of Science, Uljanov Street 46, Nizhnii Novgorod, Russia 603950

Although propadienediylidene (tricarbon, C_{3}) is a major consituent of carbon vapour and therefore a vital molecule for diagnosing the state of astrophysical and terrestial gases containing carbon, available spectroscopic data on hot C_{3} are sparse and studies limited as a consequence. As steps towards constructing a full line-list involving states with energies (⩽ 4000 cm^{−1}) we have undertaken two tasks. First, levels and corresponding line strengths have been calculated for ^{12}C^{12}C^{12}C and its isotopologues (^{12}C^{13}C^{12}C and ^{13}C^{12}C^{12}C) for 0 ⩽ J ⩽ 20 and 0 ⩽ K ⩽ 15. Partition functions and cross-sections have been deduced for temperatures ⩽ 20K and made available for absorption line studies at interstellar conditions. These results can be compared with the astronomical observations of Giessen et al. [1] Second, a MARVEL (measured vibration rotation energy levels) study has been completed: transition data covering both the ground (X^{1}Σ_{g}^{+}) and first singlet excited (A^{1}Π_{u}) states were extracted from 13 sources giving a set of 4773 transitions from which 1687 energy levels covering both the X and A state were extracted. The recent work of Martin-Drumel et al. [2] provided a particularly useful starting point for this study. Results of both studies will be reported at the conference.

[1] T.F. Giesen et al., Astron. Astrophys., 633, A120 (2020).

[2] M.-A. Martin-Drumel et al., J. Molec. Spectrosc. 391, 111734 (2023).

p-number: p217  Submitted on Fri, 16 Jun 2023 07:19:51 +0200